OpenStudy (anonymous):

Blake knows that one of the solutions to x2 – 6x + 8 = 0 is x = 2. What is the other solution? x = –6 x = –3 x = 4 x = 8

OpenStudy (anonymous):

@triciaal

jimthompson5910 (jim_thompson5910):

plug in each answer choice and see which value of x makes the left side equal to 0

jimthompson5910 (jim_thompson5910):

for example, choice A, x = -6 x^2 - 6x + 8 = 0 (-6)^2 - 6(-6) + 8 = 0 ... replace every x with -6 36 + 36 + 8 = 0 80 = 0 ... this is false, so x = -6 is NOT a solution

OpenStudy (triciaal):

given that one of the solution is x = 2 then you know (x-2) is a factor

OpenStudy (anonymous):

So C Is The Answer Correct

OpenStudy (triciaal):